OpenStudy (anonymous):

how do you write y = 3x^2 - 24x + 10 in general form?

OpenStudy (anonymous):

y = 3x² - 24x + 48 + 10 - 48 = 3(x² - 8x + 16) - 38 = 3(x - 4)² - 38

OpenStudy (anonymous):

Is that the right answer?